# HexGen: Generative Inference of Large-Scale Foundation Model over Heterogeneous Decentralized Environment

### v2

@@ Abstract.
Serving generative inference of the large-scale foundation model is a crucial component of contemporary AI applications. This paper focuses on deploying such services in a heterogeneous and decentralized setting to mitigate the substantial inference costs typically associated with centralized data centers. Towards this end, we propose HexGen, a flexible distributed inference engine that uniquely supports the asymmetric partition of generative inference computations over both tensor model parallelism and pipeline parallelism and allows for effective deployment across diverse GPUs interconnected by a fully heterogeneous network. We further propose a sophisticated scheduling algorithm grounded in constrained optimization that can adaptively assign asymmetric inference computation across the GPUs to fulfill inference requests while maintaining acceptable latency levels. We conduct an extensive evaluation to verify the efficiency of HexGen by serving the state-of-the-art Llama-2 (70B) model. The results suggest that HexGen can choose to achieve up to 2.3×2.3$\times$2.3 × lower latency deadlines or tolerate up to 4×4$\times$4 × more request rates compared with the homogeneous baseline given the same budget.
Equal contributions are indicated by $*{}^{*}$.
Correspond to Binhang Yuan (biyuan@ust.hk).
1. Introduction
@@
Foundation models are distinguished by the vast scale of parameters being trained over a substantial pre-train corpus. Such extensive training enables them to be remarkably adaptable across a broad spectrum of downstream tasks (Bommasani et al., 2021). In fact, large-scale foundation models such as GPT-4 (Bubeck et al., 2023), Llama2-70B (Touvron et al., 2023), and Falcon-170B (Institute, 2023) have essentially revolutionized the way AI systems are developed and deployed, which have nourished a large number of advanced applications such as Copilot (Github, 2023) and Bing Chat (Microsoft, 2023). In such an ecosystem, serving the generative inference requests for foundation models presents a critical challenge — given the unprecedented model scale, unlike classic machine learning models, parallel inference strategies have to be leveraged to accommodate the high computational and memory demands while ensuring low-latency generative inference outcomes.
@@
The state-of-the-art inference service of the foundation model is usually hosted in a centralized data center with homogeneous high-performance GPUs, which can be very expensive in terms of the cloud service fee. The high cost of such deployment potentially limits the democratization of this great technique. Alternatively, the deployment of the foundation model inference over a heterogeneous and decentralized environment can be a promising direction to reduce the inference cost, which has not been fully explored. The heterogeneous environment for foundation model inference service can encompass a wide range of options, including more affordable cloud services (such as spot instances (Thorpe et al., 2023; Athlur et al., 2022) and serverless computing (Guo et al., 2022)) to fully decentralized platforms (Yuan et al., 2022; Borzunov et al., 2023) that leverage a diverse set of GPUs contributed by volunteers.
@@
However, deploying foundation model inference across a heterogeneous environment presents some unique challenges. Unlike traditional machine learning models, foundation model inference consists of two different phases: a prompt phase that handles a sequence of input tokens at once and a decoding phase where output tokens are generated step-by-step. Additionally, large-scale foundation models require the adoption of specialized parallel inference strategies to effectively distribute the intensive computations across multiple GPUs. The two most commonly employed approaches are tensor model parallelism and pipeline parallelism. In terms of coordinating such a complicated distributed computation, there are some fundamental challenges stemming from the heterogeneity:
@@
  • Heterogeneous GPU computation capacity. To fully leverage the economic GPU computation power, it is essential to employ a range of GPU types, each has distinct peak FLOPS, GPU device memory bandwidth, and GPU device memory constraints. However, most (if not all) distributed implementations of foundation model inference frameworks do not accommodate this GPU diversity, as they typically assume a homogeneous GPU cluster configuration. This results in an implementation that enforces complete symmetry in the distribution of the inference computation, e.g., requiring every pipeline stage to adhere to the same tensor model parallel degree, which essentially limits the performances of the distributed inference over heterogeneous GPUs.
@@
In order to overcome these challenges, we propose HexGen, a foundation model inference system that coordinates distributed inference computation over a set of GPUs with different computation capacities connected by heterogeneous connections. We provide flexible support of asymmetric parallel execution within the scope of tensor model parallelism and pipeline parallelism to accommodate the heterogeneous GPU computation capacity. We also propose a scheduling algorithm to determine an efficient allocation of inference computation under diversified connections. Concretely, we make the following contributions:
@@
2. Preliminary
2.1. Generative inference computation.
A typical generative inference task of the transformer-based foundation model consists of two stages: i) the prefill stage that takes a prompt sequence to compute the key-value cache (KV cache) for each transformer layer of the foundation model; and ii) the decoding stage which utilizes the previous KV cache to generate new tokens step-by-step and appends the new KV cache. The inference computation can be summarized below. Denote the batch size by $b$, the prompt sequence length by $s^{\text{in}}$, the output sequence length by $s^{\text{out}}$, the hidden dimension of the transformer by $H$, and the total number of transformer layers by $L$. Given the weight matrices of a transformer layer specified by $\mathbf{w}_{K}^{\kappa},\mathbf{w}_{Q}^{\kappa},\mathbf{w}_{V}^{\kappa},\mathbf{w}_{O}^{\kappa}\in\mathcal{R}^{H\times H}$, $\mathbf{w}_{1}^{\kappa}\in\mathcal{R}^{H\times 4H}$, and $\mathbf{w}_{2}^{\kappa}\in\mathcal{R}^{4H\times H}$. During the prefill phase, the input of the $\kappa$-th layer is specified by $\mathbf{x}^{\kappa}$, and key, value, query, and output of the attention layer is specified $\mathbf{x}_{K}^{\kappa},\mathbf{x}_{V}^{\kappa},\mathbf{x}_{Q}^{\kappa},\mathbf{x}_{\text{Out}}^{\kappa}\in\mathcal{R}^{b\times s\times H}$. First, the cached key, value can be computed by:
$\begin{array}[]{cc}&\mathbf{x}_{K}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{K}^{\kappa};\quad\mathbf{x}_{V}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{V}^{\kappa}\end{array}$
@@ 2. Preliminary
$\begin{array}[]{cc}&\mathbf{x}_{Q}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{Q}^{\kappa}\\ &\mathbf{x}_{\text{Out}}^{\kappa}=f_{\text{softmax}}\left(\frac{\mathbf{x}_{Q}^{\kappa}{\mathbf{x}_{K}^{\kappa}}^{T}}{\sqrt{H}}\right)\cdot\mathbf{x}_{V}^{\kappa}\cdot\mathbf{w}_{O}^{\kappa}+\mathbf{x}^{\kappa}\\ &\mathbf{x}^{\kappa+1}=f_{\text{relu}}\left(\mathbf{x}_{\text{Out}}^{\kappa}\cdot\mathbf{w}_{1}^{\kappa}\right)\cdot\mathbf{w}_{2}^{\kappa}+\mathbf{x}_{\text{Out}}^{\kappa}\end{array}$
@@
$\begin{array}[]{cc}\mathbf{x}_{K}^{\kappa}\leftarrow\text{Concat}\left(\mathbf{x}_{K}^{\kappa},\mathbf{t}^{\kappa}\cdot\mathbf{w}_{K}^{\kappa}\right);\\ \mathbf{x}_{V}^{\kappa}\leftarrow\text{Concat}\left(\mathbf{x}_{V}^{\kappa},\mathbf{t}^{\kappa}\cdot\mathbf{w}_{V}^{\kappa}\right)\end{array}$
@@
$\begin{array}[]{cc}&\mathbf{t}_{Q}^{\kappa}=\mathbf{t}^{\kappa}\cdot\mathbf{w}_{Q}^{\kappa}\\ &\mathbf{t}_{\text{Out}}^{\kappa}=f_{\text{softmax}}\left(\frac{\mathbf{t}_{Q}^{\kappa}{\mathbf{x}_{K}^{\kappa}}^{T}}{\sqrt{H}}\right)\cdot\mathbf{x}_{V}^{\kappa}\cdot\mathbf{w}_{O}^{\kappa}+\mathbf{t}^{\kappa}\\ &\mathbf{t}^{\kappa+1}=f_{\text{relu}}\left(\mathbf{t}_{\text{Out}}^{\kappa}\cdot\mathbf{w}_{1}^{\kappa}\right)\cdot\mathbf{w}_{2}^{\kappa}+\mathbf{t}_{\text{Out}}^{\kappa}\end{array}$
2.2. Parallelism in generative inference.
Given the scale of the state-of-the-art foundation models, distributed or parallel execution is necessary. Two standard parallel strategies are usually included in inference:
Pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019) partitions the foundation model into multiple stages and serves the inference computation as a pipeline, where each GPU or (group of GPUs) handles a stage. During the inference computation, the GPU(s) serving stage-($j$) needs to send the activations to the GPU(s) serving stage-($j{+}1$). For inference computation, pipeline parallelism cannot reduce the completion time for a single request since only one stage can be active. However, the communication volume included in pipeline parallelism is much less when compared with tensor model parallelism, which is beneficial for slow GPU connections.
Tensor model parallelism (Narayanan et al., 2021) partitions the inference computation at the level of transformer layers over multiple GPUs, where the weight matrices are distributed both row-wisely and column-wisely. Two AllReduce operations are required to aggregate each layer’s output activations. Tensor model parallelism splits both the data scan and computation among a tensor model parallel group, which can effectively scale out the inference computation if the connection is fast among the group. However, when the intra-group communication is not extremely fast (i.e., not by NVLink or PCIe), tensor model parallelism can perform poorly.
3. Asymmetric Parallel Support
@@
3.1. Case Study: Parallelism over Heterogeneity
@@
Figure 1. Case study of parallel strategy over heterogeneity.
3.2. HexGen Asymmetric Parallel Support
In order to implement the asymmetric parallel support, our essential change can be summarized as follows: each pipeline parallel stage can be assigned with a different number of layers and tensor model parallel degree. From the perspective of the implementation, we make the following changes:
  • When initializing the pipeline parallel communication group, HexGen will take the configured tensor model parallel degree for each stage along with their corresponding number of allocated transformer layers;
  • Each stage will select a leader GPU which introduces lowest communication latency to GPUs in the nearby stages, and initialize an independent tensor model parallel group;
  • When executing the inference through the pipeline, only the leader node in each stage (i.e., tensor model parallel group) will send the activation to the leader GPU in the next stage; once the leader GPU receives the activation, it will broadcast this activation among its tensor model parallel group to execute the tensor model parallel computation.
In terms of system implementation, we modify the latest FlashAttention (Dao, 2023) framework by integrating this new pipeline parallel design to enable this functionality.
Table 1. Modeling the generative inference cost and limit.
@@ 3.2. HexGen Asymmetric Parallel Support
4. Scheduling over Heterogeneity
@@ 4.1. Problem Formalization
(1) $\sigma^{*}=\arg\max_{\sigma\in\Sigma}$ $\quad\mathbb{E}_{\mathbf{T}\sim\mathcal{P}}\left[\textsc{SLO}\left(\textsc{C}_{\text{comm}}\left(\sigma\right)+\textsc{C}_{\text{comp}}\left(\sigma\right)\right)\right]$
@@
4.2. Optimizing the Layout of a Pipeline
@@
(2) $\quad\quad\textsc{C}^{i}_{\text{comp}}\left(\left\{\mathbf{d}_{i,\sim}\right\}\right)+\textsc{C}^{i}_{\text{comm}}\left(\left\{\mathbf{d}_{i,\sim}\right\}\right)$
@@
(3) $\textsc{DP}\left[j;\boldsymbol{\tau}\right]=\min_{\tau_{k}\cdot\mathbf{e}_{k}\subset\mathbf{d}_{i,\sim}}\left\{\textsc{DP}\left[j-1;\boldsymbol{\tau}-\tau_{k}\cdot\mathbf{e}_{k}\right]+\right.$
@@
  Input: Memorization Buffer DP,    Current stage $j=1$,    Assigned GPU set $\boldsymbol{\tau}=\emptyset$,   Unassigned GPU set $\boldsymbol{\overline{\tau}}=\mathbf{d}_{i,\sim}$ function Estimate-Pipeline-Cost$\left(\textsc{DP},\>j,\>\boldsymbol{\tau},\>\boldsymbol{\overline{\tau}}\right)$
@@
     end forend function
4.3. Searching via Genetic Algorithm

### v3

# HexGen: Generative Inference of Large Language Model over Heterogeneous Environment

@@ Abstract
Serving generative inference of the large language model is a crucial component of contemporary AI applications. This paper focuses on deploying such services in a heterogeneous and cross-datacenter setting to mitigate the substantial inference costs typically associated with a single centralized datacenter. Towards this end, we propose HexGen, a flexible distributed inference engine that uniquely supports the asymmetric partition of generative inference computations over both tensor model parallelism and pipeline parallelism and allows for effective deployment across diverse GPUs interconnected by a fully heterogeneous network. We further propose a sophisticated scheduling algorithm grounded in constrained optimization that can adaptively assign asymmetric inference computation across the GPUs to fulfill inference requests while maintaining acceptable latency levels. We conduct an extensive evaluation to verify the efficiency of HexGen by serving the state-of-the-art Llama-2 (70B) model. The results suggest that HexGen can choose to achieve up to 2.3×2.3$\times$2.3 × lower latency deadlines or tolerate up to 4×4$\times$4 × more request rates compared with the homogeneous baseline given the same budget. Our implementation is available at https://github.com/Relaxed-System-Lab/HexGen.
Machine Learning, ICML
1 Introduction
@@
Large language models are distinguished by the vast scale of parameters being trained over a substantial pre-train corpus. Such extensive training enables them to be remarkably adaptable across a broad spectrum of downstream tasks (Bommasani et al., 2021). In fact, large language models such as GPT-4 (Bubeck et al., 2023), Llama2-70B (Touvron et al., 2023), and Falcon-180B (Institute, 2023) have essentially revolutionized the way AI systems are developed and deployed, which have nourished a large number of advanced applications. In such an ecosystem, serving the generative inference requests for large language models presents a critical challenge — given the unprecedented model scale, unlike classic machine learning models, parallel inference strategies have to be leveraged to accommodate the high computational and memory demands while ensuring low-latency generative inference outcomes.
@@
The state-of-the-art inference service of the large language model is usually hosted in a single centralized data center with homogeneous high-performance GPUs, which can be very expensive in terms of the cloud service fee. The high cost of such deployment potentially limits the democratization of this great technique. Alternatively, the deployment of the large language model inference over a heterogeneous cross-datacenter environment can be a promising direction to reduce the inference cost, which has not been fully explored. The heterogeneous environment for foundation model inference service can encompass a wide range of options, including more affordable cloud services (such as spot instances (Thorpe et al., 2023; Athlur et al., 2022) and serverless computing (Guo et al., 2022)) to even fully decentralized platforms (Yuan et al., 2022; Borzunov et al., 2023) that leverage a diverse set of GPUs contributed by volunteers in an extreme setting.
@@
However, deploying large language model inference across a heterogeneous environment presents some unique challenges. Unlike traditional machine learning models, large language model inference consists of two different phases: a prompt phase that handles a sequence of input tokens at once and a decoding phase where output tokens are generated step-by-step. Additionally, large language models require the adoption of specialized parallel inference strategies to effectively distribute the intensive computations across multiple GPUs. The two most commonly employed approaches are tensor model parallelism and pipeline parallelism. In terms of coordinating such a complicated distributed computation, there are some fundamental challenges stemming from the heterogeneity:
@@
  • Heterogeneous GPU computation capacity. To fully leverage the economic GPU computation power, it is essential to employ a range of GPU types, each has distinct peak FLOPS, GPU device memory bandwidth, and GPU device memory constraints. However, most (if not all) distributed implementations of large language model inference frameworks do not accommodate this GPU diversity, as they typically assume a homogeneous GPU cluster configuration. This results in an implementation that enforces complete symmetry in the distribution of the inference computation, e.g., requiring every pipeline stage to adhere to the same tensor model parallel degree, which essentially limits the performances of the distributed inference over heterogeneous GPUs.
@@
In order to overcome these challenges, we propose HexGen, a large language model inference system that coordinates distributed inference computation over a set of GPUs with different computation capacities connected by heterogeneous connections. We provide flexible support of asymmetric parallel execution within the scope of tensor model parallelism and pipeline parallelism to accommodate the heterogeneous GPU computation capacity. We also propose a scheduling algorithm to determine an efficient allocation of inference computation under diversified connections. Concretely, we make the following contributions:
@@
2 Preliminary
Generative inference computation. A typical generative inference task of the transformer-based foundation model consists of two stages: i) the prefill stage that takes a prompt sequence to compute the key-value cache (KV cache) for each transformer layer of the large language model; and ii) the decoding stage which utilizes the previous KV cache to generate new tokens step-by-step and appends the new KV cache. The inference computation can be summarized below. Denote the batch size by $b$, the prompt sequence length by $s^{\text{in}}$, the output sequence length by $s^{\text{out}}$, the hidden dimension of the transformer by $H$, and the total number of transformer layers by $L$. Given the weight matrices of a transformer layer specified by $\mathbf{w}_{K}^{\kappa},\mathbf{w}_{Q}^{\kappa},\mathbf{w}_{V}^{\kappa},\mathbf{w}_{O}^{\kappa}\in\mathcal{R}^{H\times H}$, $\mathbf{w}_{1}^{\kappa}\in\mathcal{R}^{H\times 4H}$, and $\mathbf{w}_{2}^{\kappa}\in\mathcal{R}^{4H\times H}$. During the prefill phase, the input of the $\kappa$-th layer is specified by $\mathbf{x}^{\kappa}$, and key, value, query, and output of the attention layer is specified $\mathbf{x}_{K}^{\kappa},\mathbf{x}_{V}^{\kappa},\mathbf{x}_{Q}^{\kappa},\mathbf{x}_{\text{Out}}^{\kappa}\in\mathcal{R}^{b\times s\times H}$. First, the cached key, value can be computed by:
$\small\begin{array}[]{cc}&\mathbf{x}_{K}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{K}^{\kappa};\quad\mathbf{x}_{V}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{V}^{\kappa}\end{array}\vspace{-0.5em}$
@@ 2 Preliminary
$\small\begin{array}[]{cc}&\mathbf{x}_{Q}^{\kappa}=\mathbf{x}^{\kappa}\cdot\mathbf{w}_{Q}^{\kappa}\\ &\mathbf{x}_{\text{Out}}^{\kappa}=f_{\text{softmax}}\left(\frac{\mathbf{x}_{Q}^{\kappa}{\mathbf{x}_{K}^{\kappa}}^{T}}{\sqrt{H}}\right)\cdot\mathbf{x}_{V}^{\kappa}\cdot\mathbf{w}_{O}^{\kappa}+\mathbf{x}^{\kappa}\\ &\mathbf{x}^{\kappa+1}=f_{\text{relu}}\left(\mathbf{x}_{\text{Out}}^{\kappa}\cdot\mathbf{w}_{1}^{\kappa}\right)\cdot\mathbf{w}_{2}^{\kappa}+\mathbf{x}_{\text{Out}}^{\kappa}\end{array}\vspace{-0.5em}$
@@
$\small\mathbf{x}_{K}^{\kappa}\leftarrow\text{Concat}\left(\mathbf{x}_{K}^{\kappa},\mathbf{t}^{\kappa}\cdot\mathbf{w}_{K}^{\kappa}\right);\>\mathbf{x}_{V}^{\kappa}\leftarrow\text{Concat}\left(\mathbf{x}_{V}^{\kappa},\mathbf{t}^{\kappa}\cdot\mathbf{w}_{V}^{\kappa}\right)\vspace{-0.5em}$
@@
$\small\begin{array}[]{cc}&\mathbf{t}_{Q}^{\kappa}=\mathbf{t}^{\kappa}\cdot\mathbf{w}_{Q}^{\kappa}\\ &\mathbf{t}_{\text{Out}}^{\kappa}=f_{\text{softmax}}\left(\frac{\mathbf{t}_{Q}^{\kappa}{\mathbf{x}_{K}^{\kappa}}^{T}}{\sqrt{H}}\right)\cdot\mathbf{x}_{V}^{\kappa}\cdot\mathbf{w}_{O}^{\kappa}+\mathbf{t}^{\kappa}\\ &\mathbf{t}^{\kappa+1}=f_{\text{relu}}\left(\mathbf{t}_{\text{Out}}^{\kappa}\cdot\mathbf{w}_{1}^{\kappa}\right)\cdot\mathbf{w}_{2}^{\kappa}+\mathbf{t}_{\text{Out}}^{\kappa}\end{array}$
Parallelism in generative inference. Given the scale of the state-of-the-art large language models, distributed or parallel execution is necessary. Two standard parallel strategies are usually included in inference: Pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019) partitions the foundation model into multiple stages and serves the inference computation as a pipeline, where each GPU or (group of GPUs) handles a stage. During the inference computation, the GPU(s) serving stage-($j$) needs to send the activations to the GPU(s) serving stage-($j{+}1$). For inference computation, pipeline parallelism cannot reduce the completion time for a single request since only one stage can be active. However, the communication volume included in pipeline parallelism is much less when compared with tensor model parallelism, which is beneficial for slow GPU connections. Tensor model parallelism (Narayanan et al., 2021) partitions the inference computation at the level of transformer layers over multiple GPUs, where the weight matrices are distributed both row-wisely and column-wisely. Two AllReduce operations are required to aggregate each layer’s output activations. Tensor model parallelism splits both the data scan and computation among a tensor model parallel group, which can effectively scale out the inference computation if the connection is fast among the group. However, when the intra-group communication is not extremely fast (i.e., not by NVLink or PCIe), tensor model parallelism can perform poorly.
3 Asymmetric Parallel Implementation
@@
3.1 Case Study: Parallelism over Heterogeneity
@@
Figure 1: Case study of parallel strategy over heterogeneity.
3.2 HexGen Asymmetric Parallel Implementation
In order to implement the asymmetric parallel support, our essential change can be summarized as follows: each pipeline parallel stage can be assigned with a different number of layers and tensor model parallel degree. From the perspective of the implementation, we make the following changes: i) when initializing the pipeline parallel communication group, HexGen will take the configured tensor model parallel degree for each stage along with their corresponding number of allocated transformer layers; ii) each stage will select a leader GPU which introduces lowest communication latency to GPUs in the nearby stages, and initialize an independent tensor model parallel group; iii) when executing the inference through the pipeline, only the leader GPU in each stage (i.e., tensor model parallel group) will send the activation to the leader GPU in the next stage; once the leader GPU receives the activation, it will broadcast this activation among its tensor model parallel group to execute the tensor model parallel computation. In terms of system implementation, we modify the latest FlashAttention (Dao, 2023) framework by integrating this new pipeline parallel design to enable this functionality.
Table 1: Modeling the generative inference cost and limit.
@@ 3.2 HexGen Asymmetric Parallel Implementation
4 Scheduling over Heterogeneity
@@ 4.1 Problem Formalization
$\sigma^{*}=\arg\max_{\sigma\in\Sigma}$ $\quad\mathbb{E}_{\mathbf{T}\sim\mathcal{P}}\left[\textsc{SLO}\left(\textsc{C}_{\text{comm}}\left(\sigma\right)+\textsc{C}_{\text{comp}}\left(\sigma\right)\right)\right]$ (1)
@@
4.2 Optimizing the Layout of a Pipeline
@@
$\quad\quad\textsc{C}^{i}_{\text{comp}}\left(\left\{\mathbf{d}_{i,\sim}\right\}\right)+\textsc{C}^{i}_{\text{comm}}\left(\left\{\mathbf{d}_{i,\sim}\right\}\right)$ (2)
@@
$\textsc{DP}\left[j;\boldsymbol{\tau}\right]=\min_{\tau_{k}\cdot\mathbf{e}_{k}\subset\mathbf{d}_{i,\sim}}\left\{\textsc{DP}\left[j-1;\boldsymbol{\tau}-\tau_{k}\cdot\mathbf{e}_{k}\right]+\right.$ (3)
@@
  Input: Memorization Buffer DP,    Current stage $j=1$,    Assigned GPU set $\boldsymbol{\tau}=\emptyset$,   Unassigned GPU set $\boldsymbol{\overline{\tau}}=\mathbf{d}_{i,\sim}$
  function Estimate-Pipeline-Cost$\left(\textsc{DP},\>j,\>\boldsymbol{\tau},\>\boldsymbol{\overline{\tau}}\right)$
@@
  end for
  end function
4.3 Searching via Genetic Algorithm